OpenStudy (anonymous):

How do u solve this equation by substitution 2x + 4y = 6 and 7x = 4 + 3y? Please help n explain

OpenStudy (zzr0ck3r):

solve either equation for x or y, then plug that into the other equation

OpenStudy (anonymous):

How?

OpenStudy (zzr0ck3r):

solve 2x+4y=6 for y y on one side and everything else on the other

OpenStudy (anonymous):

Hmm...k I think im getting wat ur saying

OpenStudy (zzr0ck3r):

subtract 2x from both sides then divide both sides by 4

OpenStudy (anonymous):

K now no

OpenStudy (zzr0ck3r):

what do you have?

OpenStudy (anonymous):

Show me how please

OpenStudy (anonymous):

Im not getting it

OpenStudy (zzr0ck3r):

2x+4y=6 subtract 4y from both sides

OpenStudy (anonymous):

Are u sure of what ur doing?

OpenStudy (anonymous):

@DebbieG please help me out here

OpenStudy (zzr0ck3r):

yes im sure, and people are not going to give you the answer 2x+4y=6 subtract 4y from both sides if you cant do this then you should go back and learn what you missed

OpenStudy (anonymous):

-4y = 2x - 6

OpenStudy (anonymous):

I guess

OpenStudy (zzr0ck3r):

\[2x+4y=6\] subtract 4y from both sides \[2x+4y\color{blue}{-4y}=6\color{blue}{-4y}\\2x+\cancel{4y}-\cancel{4y}=6-4y\\so\\2x=6-4y\] with me?

OpenStudy (zzr0ck3r):

@hubbabubba do you understand that?

OpenStudy (anonymous):

Yes I got it now thx

OpenStudy (zzr0ck3r):

well this is only the first step, there is more to do now divide each term of 2x=6-4y by 2

OpenStudy (anonymous):

Ok so its x = 3 - 2y right

OpenStudy (zzr0ck3r):

correct, now we plug this in for x in the second equation \[7\color{blue}{x} = 4 + 3y\]becomes\[7(\color{blue}{3-2y})=4+3y\]can you solve for y?

OpenStudy (anonymous):

Yea already did n I got y as 1 idk if its right tho

OpenStudy (zzr0ck3r):

yes that is correct so y=1 now we use any equation to find out what x is i would use this one x=3-2y x=3-2*1 x=1 so the solution is (1,1)

OpenStudy (anonymous):

Yaaay!!!! Im right yea I got that thanx sooo much. Thats the way I like explanation to be :)

OpenStudy (anonymous):

Ok I have another question but instead we are gonna use elimination

OpenStudy (anonymous):

Think u can help me out??

OpenStudy (anonymous):

m - n = -9 7m + 2n = 9 Use elimination method

OpenStudy (anonymous):

Help here

OpenStudy (anonymous):

Hello...anybody there?????